An electron spends about 10^-8s in an excited state before it drops to a lower state by emiting a photon. How many revolution does an electron in an n=2 bohr orbit make in 1.00*10^-8s? (ans_8.23*10^6 revolution)

Grade:12

1 Answers

vikas askiitian expert
13 years ago

velocity in any orbit = (2.18*106)/n m/s

 V2 = (2.18*106)/2 m/s

radius of any orbit =  (52.9)*10-12n2 m

r2 = (52.9*10-12)4 m

time in which electron completes one revolution is

T = 2pir2/V2

T =12*10-16                    

let it completes n crevolution so total time taken = n(12*10-16)

 n(12*10-16) = 10-8

    n = 8.23 * 106 revolutions